The heat is flowing through two cylindrical rods of same material.the diametres of the rods are in the ratio 1:2 and their lengths are in the ratio 2:1.if the temperature difference between their ends is the same,the ratio of rate of flow of heat through them will be:

Grade:12

2 Answers

askIITianexpert IITDelhi
14 years ago

Heat conduction(rate of flow of heat) is proportional to area/length everything else being identical.

So required ratio =(1/2)2 / (2/1) = 1/8

Kushagra Madhukar
3 years ago
Dear student,
Please find the answer to your question.

We know, H = KA(T2 – T1)/L
Hence, H ∝ A/L
Hence, H2/H1 = A2L1/A1L2 =(A2/A1)/(L2/L1) = (1/2)2/(2/1) [since, A ∝ d2 ]
Hence, the ratio of heat transfer = 1 : 8

Thanks and regards,
Kushagra